OpenStudy (anonymous):

bob bought popcorn hot dog soda for 9.86 popcorn is 1 dollar more than a hot dog soda costs 0.89 cents less than a hotdog

OpenStudy (aripotta):

are we trying to find the cost of each?

OpenStudy (anonymous):

yes

OpenStudy (aripotta):

p + h + s = 9.86 p = 1 + h s = h - 0.89 does that make sense so far?

OpenStudy (anonymous):

its mainly just the first 1 i need cause i can figure out the others

OpenStudy (aripotta):

p = popcorn h = hot dog s = soda

OpenStudy (anonymous):

mainly just the hotdog i need to find cause the other 2 are easy from there

OpenStudy (aripotta):

p + h + s = 9.86 p = 1 + h s = h - 0.89 (1 + h) + h + (h - 0.89) = 9.86

OpenStudy (aripotta):

(1 + h) + h + (h - 0.89) = 9.86 3h + 0.11 = 9.86 3h = 9.75 h = 3.25

OpenStudy (aripotta):

does that make sense?

OpenStudy (anonymous):

yep thanks that question is finished now

OpenStudy (aripotta):

yup

OpenStudy (anonymous):

a cell phone company charges 43.55 each month and this includes 500 mins for mins used beyond the first 500 its 0.25 per minute jims bill came to 88.30 how many minutes over 500 did he use

OpenStudy (aripotta):

43.55 + .25x = 88.30 can you solve for x now?

OpenStudy (anonymous):

i will try

OpenStudy (anonymous):

179 mins?

OpenStudy (aripotta):

that's what i got :)

OpenStudy (anonymous):

thaynes credit card bill was 709.20 this includes his charges and intrest if the new charges are 554.20 more than the intrest find the amount in charges and the amount in intrest